OpenStudy (anonymous):

find the values of x in the interval [0,2π] |tanx|=1

OpenStudy (abb0t):

You can also think of it as: \(\left| \frac{ \sin(x) }{ \cos(x) } \right|\) if it makes it easier

OpenStudy (abb0t):

I have to go, so it's x = \(\frac{ \pi }{ 4 }\) and x = \(-\frac{ \pi }{ 4 }\)

OpenStudy (anonymous):

how it comes π/4

OpenStudy (anonymous):

\[tanx=1 \therefore x= \frac{ \pi }{ 4 } \] in radians, or 45degrees Test it out on your calculator tan 45

OpenStudy (anonymous):

if it is given that

OpenStudy (anonymous):

\[0\le x \le 2\pi\] then the answer would be

OpenStudy (anonymous):

\[\frac{ \pi }{ 4 } and \frac{ 5\pi }{ 4 }\] or 45 degrees and 225 degrees

OpenStudy (anonymous):

How oit comes 5π/4

OpenStudy (anonymous):

is that absolute of tanx=1?

OpenStudy (anonymous):

If so \[tanx=\pm1\] since absolute values are always positive regardless of what answer the function provides

OpenStudy (anonymous):

between 0 and 2pi is the same as between 0 and 360degrees when you convert radians into degrees

OpenStudy (anonymous):

first of all just find for what values tanx=1 from this we must find the value for which tanx=1/1 for 45degrees opposite side=1 adjacent side=1 therefore tan45=1

OpenStudy (anonymous):

here is the part where positive and negative plays its role since both positive and negative 1 are both answers for \[\left| tanx \right|=1\] we must find the angles for which tanx=+or-1 on the unit circle

OpenStudy (anonymous):

 Each function is positive in their respective quadrants, with the exception of the first quadrant where every function has a positive outcome. If we were to find tanx=-1 then we would find the angle for the 2nd and 4th quadrant given by 180-45 and 360-45 where the angle is 45 since tan45=1 By doing this you will find the angle for which tanx=-1 which is 135degrees and 315degress

OpenStudy (anonymous):

But because we are finding \[tanx=\pm1\] we must find angles for which tan is both 1 and -1, we've already found the angles for -1 now we just have to find the angles for tan=1

OpenStudy (anonymous):

using the unit circle we find the angles for tanx=1 in the first and 3rd quadrant as such the angles would the 45 and 180+45=225